# Toward First-Principles Multi-Messenger Predictions: Coupling Nuclear Networks with GR Radiation-MHD in Gmunu

### v1

@@ Abstract
software: The results of this work were produced by utilising Gmunu (Cheong et al., 2020, 2021) and the energy-dependent M1 neutrino transport module (Cheong et al., 2023). The tabulated neutrino interaction were provided with NuLib (O’Connor, 2015). The EoS in the non-NSE region is provided with helmeos (Timmes & Swesty, 2000). The nuclear networks used in this work (i.e. the aprox13(Timmes, 1999; Timmes et al., 2000) are publicly available at https://cococubed.com/. The data of the simulations were post-processed and visualised with yt (Turk et al., 2011), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), SciPy (Virtanen et al., 2020) and Matplotlib (Hunter, 2007; Caswell et al., 2023).
@@ 1 Introduction
Significant progress has been made in incorporating reduced nuclear burning into hydrodynamics codes (Nakamura et al., 2014; Couch et al., 2015; Leung et al., 2015; Bruenn et al., 2016; Harris et al., 2017; Wongwathanarat et al., 2017; Sandoval et al., 2021; Bruenn et al., 2020; Navó et al., 2023), though these studies are restricted to Newtonian/special-relativistic frameworks. In parallel, large reaction networks such as SkyNet (Lippuner & Roberts, 2017) and WinNet (Reichert et al., 2023) are often employed in post-processing to calculate detailed nucleosynthesis from tracer particles. While this strategy enables the use of thousands of isotopes, it cannot account for the dynamical feedback of nuclear energy release during the explosion. Fully general-relativistic approaches that self-consistently combine energy-dependent radiation transport, magnetic fields, and nuclear burning remain rare. Some progress has been achieved in general relativity (Montero et al., 2012; Uchida et al., 2017; Fujibayashi et al., 2025), but existing implementations have so far been limited to simplified reaction networks, such as the CNO cycle, thereby constraining their applicability across diverse astrophysical scenarios. The present work therefore represents not only an initial step toward a more comprehensive and physically complete treatment of stellar explosions, but also a bridge between large post-processing networks and reduced in-situ networks within a fully relativistic framework—laying the foundation for future multi-messenger predictions in which nucleosynthesis feedback plays a central role.
In this work, we extend the Gmunu code (Cheong et al., 2020, 2021, 2023) by implementing composition evolution, a realistic stellar equation of state, and approximate nuclear networks. Since nuclear burning introduces stiff source terms, we evolve the reactions implicitly. As in our previous work (Cheong et al., 2022, 2023), implicit–explicit (IMEX) Runge–Kutta schemes (see, e.g., Ascher et al. 1997; Pareschi & Russo 2005; Cavaglieri & Bewley 2015) are employed to maintain stability while retaining reasonable step sizes. We validate the implementation through a hierarchy of tests, beginning with the stellar equation of state, followed by one-zone nuclear burning, and culminating in Newtonian verification tests, before demonstrating astrophysical applications in CCSNe.
The paper is organized as follows. Section 2 outlines the methodology and implementation of the stellar equation of state and nuclear networks. Section 3 presents verification tests. Section 4 demonstrates CCSNe simulations with nuclear burning. Finally, Section 5 summarizes our results and discusses future directions.
@@
2 Methods
2.1 Formulations
Gmunu adopts the $3+1$ reference-metric formalism (Montero et al., 2014; Mewes et al., 2020; Baumgarte & Shapiro, 2020). In this formalism, the metric can be written as
@@ 2.1 Formulations
2.2 Equations of State
Equations of state (EoSs) that depend on isotopic composition are necessary for nuclear hydrodynamical simulations. In this work, the Helmholtz-type stellar EoS is provided by coupling to the EoS driver helmeos (Timmes & Swesty, 2000). In this EoS, the fluid is composed of ideal gas ions, radiation, and degenerate or relativistic electrons, in local thermodynamic equilibrium. Additionally, we include the average mass excess per baryon contribution to the internal energy density:
@@ 2.2 Equations of State
is the mass excess of the $l$-th species with atomic mass $m_{l}$. This formulation allows the internal energy to remain unchanged during nuclear reactions, with the thermal energy adjusting accordingly when the composition changes (Perego et al., 2015; Hasenour & Duffell, 2025).
As noted in O’Connor & Ott (2010), the zero point for the specific internal energy $\epsilon$ is not arbitrary, as it affects the energy-momentum tensor and, thus, spacetime evolution. Rest-mass contributions must not be included in $\epsilon$, and the reference rest mass in the EoS must be consistent across regimes.
@@
2.2.1 Transition to Nuclear Statistical Equilibrium region
@@
Connecting the nuclear and stellar EoSs smoothly is desired for applications that involve both high and low density regions. The implementation of the transition between these regions is problem dependent. For example, various approaches have been proposed in the context of core-collapse supernovae (Rampp & Janka, 2002; Buras et al., 2006; Perego et al., 2015; Bruenn et al., 2020; Navó et al., 2023).
@@
In the nuclear EoS region (i.e. $T\geq T_{\rm hi}$), only $Y_{\rm e}$ is required and $X_{l}$ is not needed. In the stellar EoS region (i.e. $T<T_{\rm lo}$), the mass fractions $X_{l}$ are needed, and the electron fraction is overwritten as $Y_{\rm e}=\sum_{l}X_{l}Z_{l}/A_{l}$. In the mixed region, both $Y_{\rm e}$ and $X_{l}$ must be consistent. Here, we assume NSE compositions, where $X_{l}$ is given by an NSE solver (see Section 2.3.2). To minimize calls to the NSE solver, we evolve both $Y_{\rm e}$ and $X_{l}$ throughout the domain, monitoring the difference between the advected $Y_{\rm e}$ and $Y_{\rm e}(\{X_{l}\})$ in the mixed region. If the absolute difference exceeds $10^{-3}$, the NSE solver overwrites $X_{l}$.
Usually, despite different physics considered in nuclear and stellar EoSs, the physical properties at densities $\rho\sim 10^{7}-10^{8}~{\rm g\cdot cm^{-3}}$ are similar (Rampp & Janka, 2002). The EoSs start to differ at $\rho\gtrsim 10^{7}~{\rm g\cdot cm^{-3}}$. Moreover, the specific energy $\epsilon$ in the nuclear EoS is usually insensitive to temperature when $T\lesssim 1~{\rm MeV}\approx 11.6~{\rm GK}$. Therefore, when solving for $T$ given $\epsilon$, a solution may exist in both EoSs for $\rho\gtrsim 10^{7}~{\rm g\cdot cm^{-3}}$ and $T\lesssim 10^{10}~{\rm K}$. In the context of CCSNe, the temperature in high-density regions is usually high enough to assume NSE (Rampp & Janka, 2002; Navó et al., 2023). In this work, we prioritize the solution available in the nuclear EoS, using the stellar EoS only when necessary. We remind the reader that the implementation of EoS bridging is problem dependent; careful choice of EoSs and bridging methods is required.
2.3 Nuclear Network
@@
2.3.1 Coupling to Hydrodynamics
The timescale of the change rate of certain isotopes in the source term (7) can be comparable to or even shorter than the hydrodynamical timescale. In such cases, the reaction-induced source terms can be very stiff from a numerical point of view. Applying explicit time integration would be inefficient due to extremely strict constraints on the time step, potentially requiring prohibitively small steps to maintain stability. Therefore, nuclear reactions are often treated implicitly, for example using operator splitting (most notably Strang (Strang, 1968)(Hasenour & Duffell, 2025), first-order implicit Euler schemes (Lippuner & Roberts, 2017; Navó et al., 2023), or higher-order methods based on spectral deferred correction (Zingale et al., 2019).
In this work, we adopt implicit-explicit (IMEX) Runge-Kutta schemes (Ascher et al., 1997; Pareschi & Russo, 2005), which have been applied and tested previously in Gmunu for resistive magnetohydrodynamics (Cheong et al., 2022) and radiation transfer (Cheong et al., 2023). The key advantage of IMEX schemes is that they allow the stiff source terms from nuclear reactions to be treated implicitly, while the non-stiff hydrodynamical fluxes are handled explicitly. This approach provides stability for stiff reactions without sacrificing efficiency in the advection-dominated hydrodynamics. For implementation details of IMEX in Gmunu, we refer readers to Cheong et al. (2022).
@@ 2.3.1 Coupling to Hydrodynamics
Traditionally, nuclear reactions are treated as sets of first-order ordinary differential equations (ODEs) and solved with standard ODE integrators (Mueller, 1986; Hix & Thielemann, 1999; Timmes, 1999; Hasenour & Duffell, 2025). Here, we solve the non-linear system using a multidimensional Broyden method, falling back to the Newton-Raphson method if the Broyden method fails. The Jacobian $\partial f_{i}/\partial q_{j}$ is computed analytically, with derivatives of the mass fraction change rates provided by the nuclear network modules. The implementation of the Broyden solver and Jacobian computation follows Press et al. (1996).
@@
Shocks are typically under-resolved in astrophysical simulations due to the finite numerical resolution. To mitigate spurious results, it is common to suppress nuclear burning inside numerical shocks (Fryxell et al., 1989), ensuring that burning only occurs once physically appropriate post-shock conditions are established. In our implementation, we adopt a simplified version of the approach described in Zingale et al. (2024); Brady & Zingale (2025), tagging a cell as shocked when
@@
where $f_{\rm shock}=2/3$ by default. Here, $\nabla P$ and $\nabla\cdot\vec{v}$ are computed using slope-limited reconstructions (e.g., TVD or PPM limiters) to suppress oscillations near discontinuities and return a monotonic, cell-centred slope. This criterion effectively identifies cells with significant pressure jumps along the flow direction that are also undergoing compression, marking them as part of the shock region. Although we have not conducted an extensive parameter study ourselves, previous investigations indicate that variations of $f_{\rm shock}$ within the range 0.50.51.01.0 do not qualitatively alter the dynamics (Brady & Zingale, 2025). Alternative formulations, such as normalizing the pressure gradient by the local grid spacing rather than projecting along the velocity, have been shown to perform equivalently (Gronow et al., 2020).
@@ 2.3.2 NSE solver
with $N_{l}=A_{l}-Z_{l}$, $Q_{l}$ the nuclear binding energy, $h$ Planck’s constant, $k_{\rm B}$ Boltzmann’s constant. $g_{l}$ is the nuclear partition functions of nucleus $l$, which is in general temperature dependent. Nevertheless, in this work, we simply set $g_{l}=2$, and the inclusion of tabulating more realistic tables (e.g. Rauscher & Thielemann (2000)) is left as future work. $\mu_{l}^{\rm C}$ is the Coulomb corrections, which is obtained by using the fitting form by Chabrier & Potekhin (1998):
@@
3 Code tests
Although our ultimate goal is to apply nuclear burning in general-relativistic simulations, it is essential to benchmark the EoS and nuclear-reaction modules individually. Comparisons with well-established Newtonian test problems provide robust validation of the implementation. In this section we first test the modules and the code in the Newtonian regime, then present general-relativistic examples in Section 4.
3.1 Specific energy to temperature inversion
@@
Figure 1 shows the relative errors of temperature $T$ and specific energy $\epsilon$ for the round-trip tests with different electron fractions $Y_{\rm e}$. As discussed in Section 2.2.1, the solution $T$ of $\epsilon=\epsilon(T)$ may exist in both EoSs in regions where $\rho\gtrsim 10^{7}~{\rm g\cdot cm^{-3}}$ and $T\lesssim 10^{10}~{\rm K}$. Since the code prioritizes the solution available in the nuclear EoS, the temperature is not exactly recovered in these regions, which corresponds to the white areas in the lower-right part of the figures. Nevertheless, the relative error of the specific energy in these regions remains below $10^{-14}$ (bottom panels). Therefore, as long as no microphysics explicitly requires the temperature in these regions, the hydrodynamical evolution is expected to remain accurate.
Figure 1: Relative errors of temperature $T$ (top) and specific energy $\epsilon$ (bottom) from the round-trip tests for five electron fractions $Y_{\rm e}$. The upper-right regions enclosed by the white dashed lines denote where the nuclear EoS is applied, while the stellar or transitional (mixed) EoS is used elsewhere. The mixed regime corresponds to $5<T<5.8~{\rm GK}$ where both the stellar and nuclear EoS tables overlap in density. At densities $\rho\gtrsim 10^{7}~{\rm g~cm^{-3}}$ and temperatures $T\lesssim 10^{10}~{\rm K}$, the temperature recovery is imperfect because the code prioritizes solutions from the nuclear EoS (see Section 2.2.1). Nevertheless, the relative error in the specific energy $\epsilon$, which serves as a primitive variable in hydrodynamical evolution, remains below $10^{-14}$ across the explored parameter space.
3.2 Conserved-to-primitive conversion
Recovery of primitive variables from conservative variables in GRMHD is a non-trivial process, requiring the numerical solution of non-linear equations. As discussed in Section 2.2, Gmunu adopts the conserved-to-primitive conversion scheme of Kastaun et al. (2021). Although this scheme has been successfully applied with tabulated nuclear EoSs, it warrants re-testing due to the special bridging between stellar and nuclear EoSs.

### v2

@@ Abstract
software: The results of this work were produced by utilising Gmunu (Cheong et al., 2020, 2021) and the energy-dependent M1 neutrino transport module (Cheong et al., 2023). The tabulated neutrino interaction were provided with NuLib (O’Connor, 2015). The EoS in the non-NSE region is provided with helmeos (Timmes and Swesty, 2000). The nuclear networks used in this work (i.e. the aprox13(Timmes, 1999; Timmes et al., 2000) are publicly available at https://cococubed.com/. The data of the simulations were post-processed and visualised with yt (Turk et al., 2011), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; McKinney, 2010), SciPy (Virtanen et al., 2020) and Matplotlib (Hunter, 2007; Caswell et al., 2023).
@@ I Introduction
Significant progress has been made in incorporating reduced nuclear burning into hydrodynamics codes (Nakamura et al., 2014; Couch et al., 2015; Leung et al., 2015; Bruenn et al., 2016; Harris et al., 2017; Wongwathanarat et al., 2017; Sandoval et al., 2021; Bruenn et al., 2020; Navó et al., 2023), though these studies are restricted to Newtonian/special-relativistic frameworks. In parallel, large reaction networks such as SkyNet (Lippuner and Roberts, 2017) and WinNet (Reichert et al., 2023) are often employed in post-processing to calculate detailed nucleosynthesis from tracer particles. While this strategy enables the use of thousands of isotopes, it cannot account for the dynamical feedback of nuclear energy release during the explosion. Fully general-relativistic approaches that self-consistently combine energy-dependent radiation transport, magnetic fields, and nuclear burning remain rare. Some progress has been achieved in general relativity (Montero et al., 2012; Uchida et al., 2017, 2019; Fujibayashi et al., 2025), but most existing implementations rely on highly simplified reaction networks (e.g., CNO cycles) or lack simultaneous support for nuclear equations of state and neutrino transport, thereby constraining their applicability across diverse astrophysical scenarios. The present work therefore represents not only an initial step toward a more comprehensive and physically complete treatment of stellar explosions, but also a bridge between large post-processing networks and reduced in-situ networks within a fully relativistic framework—laying the foundation for future multi-messenger predictions in which nucleosynthesis feedback plays a central role.
In this work, we extend the Gmunu code (Cheong et al., 2020, 2021, 2023) by implementing composition evolution, a realistic stellar equation of state, and approximate nuclear networks. Since nuclear burning introduces stiff source terms, we evolve the reactions implicitly. As in our previous work (Cheong et al., 2022, 2023), implicit–explicit (IMEX) Runge–Kutta schemes (see, e.g., Ascher et al. 1997; Pareschi and Russo 2005; Cavaglieri and Bewley 2015) are employed to maintain stability while retaining reasonable step sizes. We validate the implementation through a hierarchy of tests, beginning with the stellar equation of state, followed by one-zone nuclear burning, and culminating in Newtonian verification tests, before demonstrating astrophysical applications in CCSNe.
The paper is organized as follows. Section II outlines the methodology and implementation of the stellar equation of state and nuclear networks. Section III presents verification tests. Section IV demonstrates CCSNe simulations with nuclear burning. Finally, Section V summarizes our results and discusses future directions.
@@
II Methods
II.1 Formulations
Gmunu adopts the $3+1$ reference-metric formalism (Montero et al., 2014; Mewes et al., 2020; Baumgarte and Shapiro, 2020). In this formalism, the metric can be written as
@@ II.1 Formulations
II.2 Equations of State
Equations of state (EoSs) that depend on isotopic composition are necessary for nuclear hydrodynamical simulations. In this work, the Helmholtz-type stellar EoS is provided by coupling to the EoS driver helmeos (Timmes and Swesty, 2000). In this EoS, the fluid is composed of ideal gas ions, radiation, and degenerate or relativistic electrons, in local thermodynamic equilibrium. Additionally, we include the average mass excess per baryon contribution to the internal energy density:
@@ II.2 Equations of State
is the mass excess of the $l$-th species with atomic mass $m_{l}$. This formulation allows the internal energy to remain unchanged during nuclear reactions, with the thermal energy adjusting accordingly when the composition changes (Perego et al., 2015; Hasenour and Duffell, 2025).
As noted in O’Connor and Ott (2010), the zero point for the specific internal energy $\epsilon$ is not arbitrary, as it affects the energy-momentum tensor and, thus, spacetime evolution. Rest-mass contributions must not be included in $\epsilon$, and the reference rest mass in the EoS must be consistent across regimes.
@@
II.2.1 Transition to Nuclear Statistical Equilibrium region
@@
Connecting the nuclear and stellar EoSs smoothly is desired for applications that involve both high and low density regions. The implementation of the transition between these regions is problem dependent. For example, various approaches have been proposed in the context of core-collapse supernovae (Rampp and Janka, 2002; Buras et al., 2006; Perego et al., 2015; Bruenn et al., 2020; Navó et al., 2023).
@@
In the nuclear EoS region (i.e. $T\geq T_{\rm hi}$), only $Y_{\rm e}$ is required and $X_{l}$ is not needed. In the stellar EoS region (i.e. $T<T_{\rm lo}$), the mass fractions $X_{l}$ are needed, and the electron fraction is overwritten as $Y_{\rm e}=\sum_{l}X_{l}Z_{l}/A_{l}$. In the mixed region, both $Y_{\rm e}$ and $X_{l}$ must be consistent. Here, we assume NSE compositions, where $X_{l}$ is given by an NSE solver (see Section II.3.2). To minimize calls to the NSE solver, we evolve both $Y_{\rm e}$ and $X_{l}$ throughout the domain, monitoring the difference between the advected $Y_{\rm e}$ and $Y_{\rm e}(\{X_{l}\})$ in the mixed region. If the absolute difference exceeds $10^{-3}$, the NSE solver overwrites $X_{l}$.
Usually, despite different physics considered in nuclear and stellar EoSs, the physical properties at densities $\rho\sim 10^{7}-10^{8}~{\rm g\cdot cm^{-3}}$ are similar (Rampp and Janka, 2002). The EoSs start to differ at $\rho\gtrsim 10^{7}~{\rm g\cdot cm^{-3}}$. Moreover, the specific energy $\epsilon$ in the nuclear EoS is usually insensitive to temperature when $T\lesssim 1~{\rm MeV}\approx 11.6~{\rm GK}$. Therefore, when solving for $T$ given $\epsilon$, a solution may exist in both EoSs for $\rho\gtrsim 10^{7}~{\rm g\cdot cm^{-3}}$ and $T\lesssim 10^{10}~{\rm K}$. In the context of CCSNe, the temperature in high-density regions is usually high enough to assume NSE (Rampp and Janka, 2002; Navó et al., 2023). In this work, we prioritize the solution available in the nuclear EoS, using the stellar EoS only when necessary. We remind the reader that the implementation of EoS bridging is problem dependent; careful choice of EoSs and bridging methods is required.
II.3 Nuclear Network
@@
II.3.1 Coupling to Hydrodynamics
The timescale of the change rate of certain isotopes in the source term (7) can be comparable to or even shorter than the hydrodynamical timescale. In such cases, the reaction-induced source terms can be very stiff from a numerical point of view. Applying explicit time integration would be inefficient due to extremely strict constraints on the time step, potentially requiring prohibitively small steps to maintain stability. Therefore, nuclear reactions are often treated implicitly, for example using operator splitting (most notably Strang (Strang, 1968)(Hasenour and Duffell, 2025), first-order implicit Euler schemes (Lippuner and Roberts, 2017; Navó et al., 2023), or higher-order methods based on spectral deferred correction (Zingale et al., 2019).
In this work, we adopt implicit-explicit (IMEX) Runge-Kutta schemes (Ascher et al., 1997; Pareschi and Russo, 2005), which have been applied and tested previously in Gmunu for resistive magnetohydrodynamics (Cheong et al., 2022) and radiation transfer (Cheong et al., 2023). The key advantage of IMEX schemes is that they allow the stiff source terms from nuclear reactions to be treated implicitly, while the non-stiff hydrodynamical fluxes are handled explicitly. This approach provides stability for stiff reactions without sacrificing efficiency in the advection-dominated hydrodynamics. For implementation details of IMEX in Gmunu, we refer readers to Cheong et al. (2022).
@@ II.3.1 Coupling to Hydrodynamics
Traditionally, nuclear reactions are treated as sets of first-order ordinary differential equations (ODEs) and solved with standard ODE integrators (Mueller, 1986; Hix and Thielemann, 1999; Timmes, 1999; Hasenour and Duffell, 2025). Here, we solve the non-linear system using a multidimensional Broyden method, falling back to the Newton-Raphson method if the Broyden method fails. The Jacobian $\partial f_{i}/\partial q_{j}$ is computed analytically, with derivatives of the mass fraction change rates provided by the nuclear network modules. The implementation of the Broyden solver and Jacobian computation follows Press et al. (1996).
@@
Shocks are typically under-resolved in astrophysical simulations due to the finite numerical resolution. To mitigate spurious results, it is common to suppress nuclear burning inside numerical shocks (Fryxell et al., 1989), ensuring that burning only occurs once physically appropriate post-shock conditions are established. In our implementation, we adopt a simplified version of the approach described in Zingale et al. (2024); Brady and Zingale (2025), tagging a cell as shocked when
@@
where $f_{\rm shock}=2/3$ by default. Here, $\nabla P$ and $\nabla\cdot\vec{v}$ are computed using slope-limited reconstructions (e.g., TVD or PPM limiters) to suppress oscillations near discontinuities and return a monotonic, cell-centred slope. This criterion effectively identifies cells with significant pressure jumps along the flow direction that are also undergoing compression, marking them as part of the shock region. Although we have not conducted an extensive parameter study ourselves, previous investigations indicate that variations of $f_{\rm shock}$ within the range 0.50.51.01.0 do not qualitatively alter the dynamics (Brady and Zingale, 2025). Alternative formulations, such as normalizing the pressure gradient by the local grid spacing rather than projecting along the velocity, have been shown to perform equivalently (Gronow et al., 2020).
@@ II.3.2 NSE solver
with $N_{l}=A_{l}-Z_{l}$, $Q_{l}$ the nuclear binding energy, $h$ Planck’s constant, $k_{\rm B}$ Boltzmann’s constant. $g_{l}$ is the nuclear partition functions of nucleus $l$, which is in general temperature dependent. Nevertheless, in this work, we simply set $g_{l}=2$, and the inclusion of tabulating more realistic tables (e.g. Rauscher and Thielemann (2000)) is left as future work. $\mu_{l}^{\rm C}$ is the Coulomb corrections, which is obtained by using the fitting form by Chabrier and Potekhin (1998):
@@
III Code tests
Although our ultimate goal is to apply nuclear burning in general-relativistic simulations, it is essential to benchmark the EoS and nuclear-reaction modules individually. Comparisons with well-established Newtonian test problems provide robust validation of the implementation. In this section we first test the modules and the code in the Newtonian regime, then present general-relativistic examples in Section IV.
III.1 Specific energy to temperature inversion
@@
Figure 1 shows the relative errors of temperature $T$ and specific energy $\epsilon$ for the round-trip tests with different electron fractions $Y_{\rm e}$. As discussed in Section II.2.1, the solution $T$ of $\epsilon=\epsilon(T)$ may exist in both EoSs in regions where $\rho\gtrsim 10^{7}~{\rm g\cdot cm^{-3}}$ and $T\lesssim 10^{10}~{\rm K}$. Since the code prioritizes the solution available in the nuclear EoS, the temperature is not exactly recovered in these regions, which corresponds to the white areas in the lower-right part of the figures. Nevertheless, the relative error of the specific energy in these regions remains below $10^{-14}$ (bottom panels). Therefore, as long as no microphysics explicitly requires the temperature in these regions, the hydrodynamical evolution is expected to remain accurate.
Figure 1: Relative errors of temperature $T$ (top) and specific energy $\epsilon$ (bottom) from the round-trip tests for five electron fractions $Y_{\rm e}$. The upper-right regions enclosed by the white dashed lines denote where the nuclear EoS is applied, while the stellar or transitional (mixed) EoS is used elsewhere. The mixed regime corresponds to $5<T<5.8~{\rm GK}$ where both the stellar and nuclear EoS tables overlap in density. At densities $\rho\gtrsim 10^{7}~{\rm g~cm^{-3}}$ and temperatures $T\lesssim 10^{10}~{\rm K}$, the temperature recovery is imperfect because the code prioritizes solutions from the nuclear EoS (see Section II.2.1). Nevertheless, the relative error in the specific energy $\epsilon$, which serves as a primitive variable in hydrodynamical evolution, remains below $10^{-14}$ across the explored parameter space.
III.2 Conserved-to-primitive conversion
Recovery of primitive variables from conservative variables in GRMHD is a non-trivial process, requiring the numerical solution of non-linear equations. As discussed in Section II.2, Gmunu adopts the conserved-to-primitive conversion scheme of Kastaun et al. (2021). Although this scheme has been successfully applied with tabulated nuclear EoSs, it warrants re-testing due to the special bridging between stellar and nuclear EoSs.